OpenStudy (anonymous):

Solve using the substitution method. Show your work. If the system has no solution or an infinite number of solutions, state this. -8x – 4y = -68 x + 3y = 16

OpenStudy (anonymous):

can you solve that second equation for X ?

OpenStudy (anonymous):

first equation should be for x

OpenStudy (anonymous):

ok then, what does X equal in the first equation?

OpenStudy (anonymous):

sorry, have to go, but someone smarter than I will help you, I am sure

OpenStudy (zepp):

hai there, in this question, you better get the shortest equation which has only 1 variable on 1 side

OpenStudy (anonymous):

solving by itself i got 0

OpenStudy (zepp):

what gives 0?

OpenStudy (zepp):

There: x + 3y = 16 This one is the shortest equation, and x seems to be the perfect variable to be forever alone :) x = 16- 3y

OpenStudy (zepp):

Then you can subtitute this x = 16-3y in the first equation, so you only have 1 variable in the whole equation, makes it lot easier to solve.

OpenStudy (anonymous):

so its easier to use the shortest equation first

OpenStudy (zepp):

Yes, if you don't want a huge equation.

OpenStudy (anonymous):

i was trying to figure out where you would switch the equation around

OpenStudy (zepp):

Little hint there: -8x - 4y = -68 There's lot of negative there and they could be taken out. 8x + 4y = 68 is much better to look at. Now all you have to do if to replace x by (16 - 3y) and find y, once you have it, you can get the x

OpenStudy (zepp):

8x + 4y = 68 8(16-3y) + 4y = 68 128 - 24y + 4y= 68 128- 20y = 68 128 - 68 = 20y 60 =20y y = 3

OpenStudy (zepp):

With this, you can find x easily, knowing y = 3 x + 3y = 16 x + 3(3) = 16 x + 9 = 16 x = 7

OpenStudy (anonymous):

It looks really simple if i always get lost on either the second step or don't know when to flip the negative signs.

OpenStudy (zepp):

You can work without changing the signs, but it's just more difficult, when you have something like -a - b = -c, the same thing as a + b = c

OpenStudy (zepp):

There's a proof: http://puu.sh/pQzo

OpenStudy (anonymous):

I was using this formula Ax+By=C which is the same that you was using i see. May I'm I'm just being hard on myself cause I know that I know it but haven't used it since I graduated in 2001.

OpenStudy (anonymous):

I was using this formula Ax+By=C which is the same that you was using i see. May I'm I'm just being hard on myself cause I know that I know it but haven't used it since I graduated in 2001.